# SynthScribe: Deep Multimodal Tools for Synthesizer Sound Retrieval and Exploration

### 07 Dec 2023 v1

@@ Abstract.
Synthesizers are powerful tools that allow musicians to create dynamic and original sounds. Existing commercial interfaces for synthesizers typically require musicians to interact with complex low-level parameters or to manage large libraries of premade sounds. To address these challenges, we implement SynthScribe — a fullstack system that uses multimodal deep learning to let users express their intentions at a much higher level. We implement features which address a number of difficulties, namely 1) searching through existing sounds, 2) creating completely new sounds, 3) making meaningful modifications to a given sound. This is achieved with three main features: a multimodal search engine for a large library of synthesizer sounds; a user centered genetic algorithm by which completely new sounds can be created and selected given the users preferences; a sound editing support feature which highlights and gives examples for key control parameters with respect to a text or audio based query. The results of our user studies show SynthScribe is capable of reliably retrieving and modifying sounds while also affording the ability to create completely new sounds that expand a musicians creative horizon.
Sound Synthesis, Sound Editing, Multimodal Deep Learning
@@
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 1. Introduction
To this end, we present SynthScribe (Fig. 1B) — a full stack system that leverages multimodal deep learning to help musicians work with an existing synthesizer. Guided by our formative interviews with musicians, we implement three features which leverage LAION-CLAP — a multimodal deep learning model pretrained on music and general audio. We first implement a multimodal search that allows users to navigate through a bank of synthesizer sounds using either text or existing synthesizer sounds. We also create a user-centered genetic algorithm where hundreds of new synthesizer sounds can be created by mixing together a list of a musician’s favourite presets; these new sounds are then recorded and embedded with LAION-CLAP on the fly, allowing users to rapidly discover completely new soundscapes through the multimodal search feature. Once a user has found a sound that nearly meets their desires, they can make meaningful edits to it using the preset modification feature which highlights important groups of parameters with respect to a text or audio query and provides examples of how to modify those parameters to achieve a desired effect.
@@
  • A novel system, SynthScribe, that enables the use of text and audio for the intuitive control of a synthesizer.
  • The results of a user study evaluating our systems ability to retrieve and modify existing synthesizer sounds.
  • The results of free usage observations with musicians showing how musicians of various skill levels interact with SynthScribe.
@@ 4. Design Goals
D2. Supporting Preset Modification. SynthScribe should also enhance a musicians ability to meaningfully change preset sounds, enabling users to adapt a synth sound to their needs or to rapidly incorporate feedback from other stakeholders.
D3. Helping Users Go Beyond Preset Sounds. SynthScribe should also enable users to create new sounds that go beyond presets, driving even novice users towards originality.
@@ 5. SynthScribe
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 5.3. Preset Modification
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 5.4. Supporting Features
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 6.3. Genetic Mixing
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@
The mechanisms for genetic mixing can be described using the language of Genetic Algorithms (GA). Initially, the user selects a group of fit presets to be bred together in order to create a new generation of fit presets. To create this new generation of individuals, we breed pairs of preset sounds by having a child randomly inherit whole groups of parameters from one parent or another. The breeding mechanism is displayed in Fig. 6. Diva and many other synthesizers are broken down into panels where each panel contains a group of parameters to control a specific aspect of the sound. For Diva, we recognize 13 such groups. In a breeding operation between parent A and parent B, two children are always created. For a specific group of parameters, the first child will randomly inherit either parent As or parent Bs group with equal probability. The second child will always receive each group that the first child did not inherit. For each pair of parent presets, we complete 5 breeding operations resulting in a total of 10 children per pair. This operation is equivalent to a uniform crossover over the groups of parameters. Upon creating a new generation of presets, they must be recorded and embedded to maintain search functionalities. This is achieved by first using the Python library DawDreamer (Braun, 2021) to record each preset sound faster than real-time and then using LAION-CLAP to embed each of these new recordings.
@@ 6.4. Preset Modification
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@ 7.1. Participants
We recruited 8 participants in total for the study (Mean age=25.0, STD=2.9). Participants were not required to have any level of musical experience. Only two participants actively played an instrument and none had extensive experience using synthesizers. The study took place across two independent tasks which happened on different days where each task lasted approximately 30 minutes. Participants were compensated with a total of 20 CAD.
@@ 8.1. Participants
Our study consisted of 6 musicians who had experience creating music. Two were professional musicians that make music for a living (P2, P5) and the rest were hobbyists who made music for their own enjoyment. Both of the professional musicians had received formal instruction on synthesizers in an academic setting with P2 describing themselves as an expert and P5 an intermediate. The hobbyists had at most encountered synthesizers when making music and described themselves as either novices or beginners. All musicians had at least basic keyboard skills. Each were compensated with 20 CAD.
@@ 8.3. Results
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
@@
In Fig. 8 we summarize the subjective ratings of SynthScribe. Our participants broadly agreed that SynthScribe is easy to use and useful overall in addition to imparting relatively low mental demand and effort while causing little frustration. The overall results of these studies were positive. P2, a professional musician, mentioned that ”even though like I know a lot about synthesis and can probably get pretty far with just the DIVA interface, this actually saved me a lot of time.” The other professional musician, P5, mentioned that the system would be most useful for them when learning to use a new synthesizer. Participants (P1, P2, P3) also mentioned that they enjoy the discoverability aspects of the system and described situations where they arrived at fitting but surprising new sounds. We also received invaluable constructive feedback that can inspire future directions. Below, we further analyze the feedback we received on each feature.
@@ 8.4. Multimodal Search
We observed that some participants use the multimodal search in a step-wise fashion (P1, P3, P6). First, they ran a text search with a general description of their idea and then they ran an audio search on a sound they like. P1, for example, started by searching for the sound of a piccolo and then narrowed their search by running an audio search on a sound they liked which led to a recorder sound that they felt to be most useful. P3 described the advantage of a general initial text search when they were looking for ambient sounds: ”My goal was very loose and it gave me some things that, for example, the bass being a candidate for the kind of ambient sound I want, I wouldn’t have thought of that myself.” Participants sometimes had difficulties with the query format. P4, for example, thought typing ”the sound of …” every time was ”completely irrelevant.” They also suggested that it might be better to have a pop-up window containing the results of an audio search to avoid losing the original content of the initial text search.
@@ 8.5. Genetic Mixing
The Genetic Mixing feature was appreciated by participants for the pleasant but unexpected sounds it created (P1, P2, P3, P4). P4 elaborated on this by saying ”I think the mix feature allows you to kind of have these weird overlaps that you wouldn’t normally have in a synth and that was really cool.” The fact that our mixing algorithm achieved aesthetically pleasing and surprising results is bolstered by the reviews of the Multimodal Search due to the fact that users can search through new generations of sounds using the same search strategies as before. Some participants expected the mixes to be slightly more thorough. P2, for example, noted that they expected the Genetic Mixing to do some level of interpolation between the continuous parameters in the favorites list instead of just swapping parameters. P3 mentioned that a useful addition would be a system that automatically names the new synthesizer sounds as opposed to the current generic naming convention.
@@ 8.6. Preset Modification
The preset modification feature was used by participants to make quick directed changes. Both professional musicians, P2 and P5, used this feature to make the final adjustments when attempting to replicate synthesizer sounds from songs they’d chosen. To this end, P2 claimed that ”even though like I know a lot about synthesis and can probably get pretty far with just the DIVA interface, this actually saved me a lot of time. P1 noted that when using synthesizers in the past ”the search space seemed unlimited” but that this feature provides ”a really fast way to just iterate through all of the different possibilities out there”. They also claimed that the parameter group highlighter was a useful indicator of importance. Other participants highlight some non-intuitive aspects of this feature. P4 and P6 believed the numerical labeling of the examples corresponded to an increasing quantity. P3 felt that having to use nouns in the text queries (e.g. ”the sound of a harsh synthesizer”) required more mental effort than just typing an adjective or command (e.g. ”more harsh”). P4 outright stated that they disliked the feature because it felt like it required them to understand the functionality of the different parameter groups.
@@ 9.1. Facilitating Synthesizer Use with Multimodal Deep Learning
Below we list some suggestions for future work that may address some limitations of SynthScribe. SynthScribe is built on top of one synthesizer but in practice, musicians interact with many synthesizers. It would be useful to design a system with similar features but manages many synthesizers simultaneously. It could also be useful to tailor a multimodal deep-learning model to synthesizer sounds. This could include training a model on text-audio pairs including synthesizer jargon or common words that people use to describe timbres (e.g. dark, bright, mellow, harsh, etc.). It’s also important to note that people tend to describe sounds with their own idiosyncrasies which could lend this design problem to systems that leverage personalization. To highlight which parameters are important to make a desirable modification, it could be useful to collect a larger bank of synthesizer sounds which would prevent the failure case that we describe above. It could also be useful to calculate importance based on fewer examples which would prevent unrelated examples from clouding the calculation. Additionally, relevant prior work has already focused on developing intuitive sliders for synthesizers (Huang et al., 2014) but used a synthesizer with relatively few parameters and required user annotations. Future work might use the preset modification approach to help the user isolate a subset of important parameters quickly which could then be used in tandem with other techniques to provide intuitive sliders with fewer user annotations and effort.
@@ 10. Conclusion
We implemented a novel system, SynthScribe, which used multimodal deep learning as a foundation for several features designed to facilitate the use of synthesizers for musicians. The results of our user studies show that our system is capable of helping musicians search for and modify synthesizer sounds effectively while also affording them the opportunity to create new and surprising sounds that inspire creativity. We outline that that future work can aim to tailor multimodal models to the language used to describe timbres but may also attempt benefit from incorporating aspects of personalization. We highlight that our participants particularly enjoyed sounds that they liked but did not expect. We recommend that future work may also try new methods that predict sounds that musicians will enjoy but were not explicitly asked for.
@@ A.2. SynthScribe Interface
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
CAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 20 Feb 2024 v2

@@ Abstract.
Synthesizers are powerful tools that allow musicians to create dynamic and original sounds. Existing commercial interfaces for synthesizers typically require musicians to interact with complex low-level parameters or to manage large libraries of premade sounds. To address these challenges, we implement SynthScribe — a fullstack system that uses multimodal deep learning to let users express their intentions at a much higher level. We implement features which address a number of difficulties, namely 1) searching through existing sounds, 2) creating completely new sounds, and 3) making meaningful modifications to a given sound. This is achieved with three main features: a multimodal search engine for a large library of synthesizer sounds; a user centered genetic algorithm by which completely new sounds can be created and selected given the users preferences; a sound editing support feature which highlights and gives examples for key control parameters with respect to a text or audio based query. The results of our user studies show SynthScribe is capable of reliably retrieving and modifying sounds while also affording the ability to create completely new sounds that expand a musicians creative horizon.
@@
A comic strip depicting a musicians experience using a synthesizer without SynthScribe and their experience using a synthesizer with SynthScribe. Without SynthScribe, the user is frustrated with the complicated interface for the synthesizer and is unable to create the harsh sound that they desire. With SynthScribe, they are considerably more satisfied and are able to create the harsh sound they are looking for.
@@ 1. Introduction
To this end, we present SynthScribe (Fig. 1B) — a full stack system that leverages multimodal deep learning to help musicians work with an existing synthesizer. Guided by our formative interviews with musicians, we implement three features which leverage LAION-CLAP (Wu* et al., 2023) — a multimodal deep learning model pretrained on music and general audio. We first implement a multimodal search that allows users to navigate through a bank of synthesizer sounds using either text or existing synthesizer sounds. We also create a user-centered genetic algorithm where hundreds of new synthesizer sounds can be created by mixing together a list of a musician’s favourite presets; these new sounds are then recorded and embedded with LAION-CLAP on the fly, allowing users to rapidly discover completely new soundscapes through the multimodal search feature. Once a user has found a sound that nearly meets their desires, they can make meaningful edits to it using the preset modification feature which highlights important groups of parameters with respect to a text or audio query and provides examples of how to modify those parameters to achieve a desired effect.
@@
  • A novel system, SynthScribe, that uses multimodal deep learning to enable the use of text and audio for the intuitive control of a synthesizer.
  • The results of a user study evaluating our systems ability to retrieve and modify existing synthesizer sounds; thus, evaluating LAION-CLAP’s effectiveness as a foundation for these tasks.
  • The results of free usage observations with musicians demonstrating that our system assists musicians in efficiently finding and creating new sounds while also providing them with unexpected sounds that inspire new creative directions.
@@ 4. Design Goals
D2. Facilitating Finetuning of Existing Presets. SynthScribe should also enhance a musician’s ability to meaningfully change preset sounds, enabling users to adapt a synth sound to their needs or to rapidly incorporate feedback from other stakeholders.
D3. Supporting the Creation of New Sounds. SynthScribe should also enable users to create new sounds that go beyond presets, driving even novice users towards originality.
@@ 5. SynthScribe
Flowchart depicting the SynthScribe workflow. The first cell depicts the multimodal search feature which uses LAION-CLAP to search through lists of synthesizer sounds. This cell is connected to the genetic mixing cell which shows how synthesizer sounds can be selected as favourites and then mixed to create completely new sounds. The multimodal search cell is also connected to the Preset Modification Cell which takes in a synthesizer sound and modifies it with respect to a text query (e.g. ”more harsh”).
@@
A screenshot of depicting the interface for the multimodal search and genetic mixing features of SynthScribe.
@@ 5.3. Preset Modification
A screenshot of depicting the interface for the preset modification feature of SynthScribe.
@@ 5.4. Supporting Features
A screenshot depicting the Oscilloscope and synthesizer parameter display.
@@ 6.3. Genetic Mixing
A visualization that demonstrates the uniform crossover algorithm used for the genetic mixing feature.
@@
The genetic mixing feature can be described using the language of Genetic Algorithms (GA). Initially, the user selects a group of fit presets to be bred together in order to create a new generation of fit presets. To create this new generation of individuals, we breed pairs of preset sounds by having a child randomly inherit whole groups of parameters from one parent or another (Fig. 6). The parameters of Diva and many other synthesizers are broken down into groups of parameters that control specific aspects of the sound. For Diva, we recognize 13 such groups. In a breeding operation between parent A and parent B, two children are always created. For a specific group of parameters, the first child will randomly inherit either parent As or parent Bs group with equal probability. The second child will always receive each group that the first child did not inherit. For each pair of parent presets, we complete 5 breeding operations resulting in a total of 10 children per pair. This operation is equivalent to a uniform crossover over the groups of parameters. Upon creating a new generation of presets, they must be recorded and embedded to maintain search functionalities. This is achieved by first using the Python library DawDreamer (Braun, 2021) to record each preset sound faster than real-time and then using LAION-CLAP to embed each of these new recordings.
@@ 7.1. Participants
We recruited 8 participants in total for the study (Mean age=25.0, STD=2.9). Participants were not required to have any level of musical experience. Only two participants actively played an instrument and none had extensive experience using synthesizers. In this regard, we followed prior work which used inexperienced participants for discerning differences in timbre. (Huang et al., 2018). We leave engagement with musicians for our free usage observations in section 8. The study took place across two independent tasks which happened on different days where each task lasted approximately 30 minutes. Participants were compensated with a total of 20 CAD.
@@ 8.1. Participants
Our study consisted of 6 musicians who had experience creating music. Two were professionals that make music for a living (P1, and P2) and the rest were hobbyists who made music for their own enjoyment. The professionals had participated in the formative studies and there was no overlap with the participant group in subsection 7.1. Both of the professional musicians had received formal instruction on synthesizers in an academic setting with P2 describing themselves as an expert and P1 an intermediate. The hobbyists had at most encountered synthesizers when making music and described themselves as either novices or beginners. All musicians had at least basic keyboard skills. Each were compensated with 20 CAD.
@@ 8.3. Results
This displays the results of the questionnaire. Participants widely agree that each feature is useful or easy to use with one participant disagreeing that the search feature was easy to use and another disagreeing that it was easy to meaningfully modify the synthesizer sounds with SynthScribe.
@@
In Fig. 8 we summarize the subjective ratings of SynthScribe. Our participants broadly agreed that SynthScribe is easy to use and useful overall in addition to imparting relatively low mental demand and effort while causing little frustration. The overall results of these studies were positive. P2, a professional musician, mentioned that ”even though like I know a lot about synthesis and can probably get pretty far with just the DIVA interface, this actually saved me a lot of time.” The other professional musician, P1, mentioned that the system would be most useful for them when learning to use a new synthesizer. Participants (P2, P3, P5) also mentioned that they enjoy the discoverability aspects of the system and described situations where they arrived at fitting but surprising new sounds. We also received invaluable constructive feedback that can inspire future directions. Below, we further analyze the feedback we received on each feature.
@@ 8.4. Multimodal Search
We observed that some participants use the multimodal search in a step-wise fashion (P3, P5, P6). First, they ran a text search with a general description of their idea and then they ran an audio search on a sound they like. P5, for example, started by searching for the sound of a piccolo and then narrowed their search by running an audio search on a sound they liked which led to a recorder sound that they felt to be most useful. P3 described the advantage of a general initial text search when they were looking for ambient sounds: ”My goal was very loose and it gave me some things that, for example, the bass being a candidate for the kind of ambient sound I want, I wouldn’t have thought of that myself.” Participants sometimes had difficulties with the query format. P4, for example, thought typing ”the sound of …” every time was ”completely irrelevant.” They also suggested that it might be better to have a pop-up window containing the results of an audio search to avoid losing the original content of the initial text search.
@@ 8.5. Genetic Mixing
The Genetic Mixing feature was appreciated by participants for the pleasant but unexpected sounds it created (P2, P3, P4, P5). P4 elaborated on this by saying ”I think the mix feature allows you to kind of have these weird overlaps that you wouldn’t normally have in a synth and that was really cool.” The fact that our mixing algorithm achieved aesthetically pleasing and surprising results is bolstered by the reviews of the Multimodal Search due to the fact that users can search through new generations of sounds using the same search strategies as before. Some participants expected the mixes to be slightly more thorough. P2, for example, noted that they expected the Genetic Mixing to do some level of interpolation between the continuous parameters in the favorites list instead of just swapping parameters. P3 mentioned that a useful addition would be a system that automatically names the new synthesizer sounds as opposed to the current generic naming convention.
@@ 8.6. Preset Modification
The preset modification feature was used by participants to make quick directed changes. Both professional musicians, P1 and P2, used this feature to make the final adjustments when attempting to replicate synthesizer sounds from songs they’d chosen. To this end, P2 claimed that ”even though like I know a lot about synthesis and can probably get pretty far with just the DIVA interface, this actually saved me a lot of time. P5 noted that when using synthesizers in the past ”the search space seemed unlimited” but that this feature provides ”a really fast way to just iterate through all of the different possibilities out there”. They also claimed that the parameter group highlighter was a useful indicator of importance. Other participants highlight some non-intuitive aspects of this feature. P4 and P6 believed the numerical labeling of the examples corresponded to an increasing quantity. P3 felt that having to use nouns in the text queries (e.g. ”the sound of a harsh synthesizer”) required more mental effort than just typing an adjective or command (e.g. ”more harsh”). P4 outright stated that they disliked the feature because it felt like it required them to understand the functionality of the different parameter groups.
@@ 9.1. Facilitating Synthesizer Use with Multimodal Deep Learning
Below we list some suggestions for future work that may address some limitations of SynthScribe. SynthScribe is built on top of one synthesizer but in practice, musicians interact with many synthesizers. It would be useful to design a system with similar features that manages many synthesizers simultaneously. It could also be useful to tailor a multimodal deep-learning model to synthesizer sounds. This could include training a model on text-audio pairs including synthesizer jargon or common words that people use to describe timbres (e.g. dark, bright, mellow, harsh, etc.). It’s also important to note that people tend to describe sounds with their own idiosyncrasies which could lend this design problem to systems that leverage personalization. To highlight which parameters are important to make a desirable modification, it could be useful to collect a larger bank of synthesizer sounds which would prevent the failure case that we describe above. It could also be useful to calculate importance based on fewer examples which would prevent unrelated examples from clouding the calculation. Additionally, relevant prior work has already focused on developing intuitive sliders for synthesizers (Huang et al., 2014) but used a synthesizer with relatively few parameters and required user annotations. Future work might use the preset modification approach to help the user isolate a subset of important parameters quickly which could then be used in tandem with other techniques to provide intuitive sliders with fewer user annotations and effort.
@@ 10. Conclusion
We implemented a novel system, SynthScribe, which used multimodal deep learning as a foundation for several features designed to facilitate the use of synthesizers for musicians. The results of our user studies show that our system is capable of helping musicians search for and modify synthesizer sounds effectively while also affording them the opportunity to create new and surprising sounds that inspire creativity. We outline that future work can aim to tailor multimodal models to the language used to describe timbres but may also benefit from incorporating aspects of personalization. We highlight that our participants particularly enjoyed sounds that they liked but did not expect. We recommend that future work may also try new methods that predict sounds that musicians will enjoy but were not explicitly asked for.
11. Acknowledgements
We would like to acknowledge the guidance we received from various collaborators that shaped the direction of SynthScribe. Particularly Nicolas Jonason who referred us to LAION-CLAP as a candidate foundation model and Marcel Castillo who provided a breadth of musical perspective in the early stages of the project. We would also like to thank Dani Oore who remixed one of his and Andrew Staniland’s compositions by using SynthScribe to source new preset sounds  [^8] Mythos Synthscribe Remix: Deus In Machina by Dani Oore and Andrew Staniland: https://soundcloud.com/dani-oore/mythos-synthscribe-remix-deus-in-machina.
@@ A.2. SynthScribe Interface
[…]